# The Monge optimal transport barycenter problem

### v1

@@ 11.3.1 Initial determination of the $\{S^{0}_{j}\}$
    and define
    $S^{0}_{j}=\frac{S_{j}}{\beta_{*}^{2}}.$
The procedure so far is automatic and based exclusively on the datapoints $\{z_{i}\}$. Yet there are at least two reasons why we may want to add one or more free parameters to the determination of the bandwidths. One is that, as discussed above, the ideal $S_{j}$ should depend not just on the $\{z_{i}\}$ but also on their relation to the $\{x_{i}\}$. One straightforward way to address this dependence is through cross-validation over such free parameters. The second reason is that often the bandwidths are determined not by the data alone but also from the scales that one would like to resolve, as in the multi time-scale analysis of ground temperature of Section 9.6. In view of this, we divide our $S_{j}$ by an externally provided constant ${\gamma_{z}}^{2}$, which we either set based on the scales that we have chosen to resolve, as in Section 9.6, or cross-validate over, as in the prediction of global sea surface temperature in Section 9.7.
11.4 The case of time-like variables $z$
The discussion above applies to bounded variables $z$, all smooth functions of which can be approximated through kernels centered at a relatively small number of well-chosen points $\{z^{c}_{i}\}$. Yet in time series-analysis, some of the $z$’s can consist of time itself under different scalings (such as $z_{3,4}$ in the ground-temperature example of Section 9.6.) The number of centers required for such time-like variables grows linearly with the time-extent of the data, which could be very large. From a practical perspective, the problem of such secular growth is that the computation of the orthogonal matrix $Q_{z}$ from $F$ requires finding the dominant singular components of a potentially vary large matrix $F$, a computationally costly task.
Yet this problem comes with its own solution. Even though $F$ is large, it is also very sparse, since $K(t_{1},t_{2})$ can be made to vanish for time pairs such that $|t_{2}-t_{1}|$ is much larger than the kernel’s bandwidth. The computational cost of finding the principal components of large but sparse matrices grows only linearly with the number of a matrix rows.
@@ Acknowledgments
Data availability statement
Ground-level temperature data are available from the NOAA U.S. Climate Reference Network. Global sea surface temperature data can be accessed from the Met Office Hadley Centre. All other data used in this study are simulated as described in the corresponding subsections.

### v2

@@ Appendix A Appendix: a data-adapted functional space
There are at least two reasons why we may want to add one or more free parameters to the determination of the bandwidths. One is that the ideal $S_{j}$ should depend not just on the $\{z_{i}\}$ but also on their relation to the $\{x_{i}\}$, and one straightforward way to address this dependence is through cross-validation over such free parameters. The second reason is that often the bandwidths are determined not by the data alone but also from the scales that one seeks to resolve, as in the multi time-scale analysis of ground temperature of Section 9.6. In view of this, we divide $S^{0}_{j}$ by an externally provided constant ${\gamma_{z}}^{2}$, which we either cross-validate over or set based on the scales that we seek to resolve.